# DAmodel: Hierarchical Bayesian Modelling of DA White Dwarfs for Spectrophotometric Calibration

### v1

@@ 2.3 Ground-Based Optical Spectra
Ground-based spectra were obtained with the Gemini North and South telescopes, the Multiple Mirror Telescope (MMT) and SOAR, as described in Calamida et al. (2019) and Calamida et al. (2022b). Spectroscopic data were obtained using the Gemini Multi-Object Spectrograph (GMOS) at both Gemini North and South observatories (Calamida et al., 2019). Observations were conducted in queue mode, initially with a $1.5^{\prime\prime}$ slit, which was later narrowed to a $1.0^{\prime\prime}$ slit. The final Gemini spectra cover a continuous range from 3500 Å to 6360 Å with a dispersion of 0.92 Å/pixel (Calamida et al., 2019). However, it was determined that these spectra were insufficient for use as spectrophotometric standards due to several factors: instrumental limitations (such as GMOS’s limited blue throughput), the absence of standard CALSPEC stars (Bohlin et al., 2014, 2020) in the observing runs, and a lack of parallactic angle alignment during observations, resulting in atmospheric dispersion losses (Calamida et al., 2019). This led to further observations at the MMT Observatory to enhance the spectral data.
At MMT, DA WDs were observed using the Blue Channel spectrograph with a 300 line $\text{mm}^{-1}$ grating. Most observations employed a $1.0^{\prime\prime}$ slit, with one epoch using a $1.25^{\prime\prime}$ slit (Calamida et al., 2019). The MMT spectra cover a wavelength range from 3400 Å to 8400 Å at a dispersion of 1.95 Å/pixel. All observations at MMT were taken at the parallactic angle, and standard stars were observed on the same night. The spectrographs did not operate on a linear pixel-to-angstrom scale, so the data were resampled onto a uniform linear grid after extraction, with a sampling of 1 Å/pixel for GMOS and 2 Å/pixel for MMT (Calamida et al., 2019). Finally, the spectra were flux-calibrated using custom routines designed for calibrating SNe Ia spectra (Matheson et al., 2008). Flux calibration was performed using a low-order polynomial with knot spacing of several hundred angstroms, making it insensitive to the sampling scale of 1–2 Å.
Approximately 50 southern white dwarf candidates were observed using the Goodman spectrograph at SOAR ($1.07^{\prime\prime}$ slit, $1.99^{\prime\prime}$/pixel, 3850–7100 Å coverage) over several runs in 2016 and 2017. A list of candidates observed and a log of observations are in (Calamida et al., 2022b).
@@ 2.4 Space-Based STIS UV Spectra
To better characterize and ultimately correct for dust effects, the spectroscopic observations were extended into the UV regime. This expansion was critical for capturing the 2175 Å bump in the dust relation, where dust extinction is most pronounced. These UV observations aimed to identify potential "anomalous" reddening sources that deviate from current extinction relations. To accurately quantify extinction and minimize observational noise, the HST Space Telescope Imaging Spectrograph (STIS) was used for data collection. The target selection criteria included the brighter standards, with magnitudes of 15 $\leq$ F275W $\leq$ 18, and known $A_{V}$ values between 0 and 1.
Spectra for 21 WDs were obtained using the STIS FUV and NUV detectors with the G140L and G230L gratings, respectively, under program 16764 led by PI Gautham Narayan. Both FUV and NUV observations employed a $52\times 2$ arcsecond slit width, covering wavelength ranges of 1150 to 1700 Å for FUV and 1570 to 3180 Å for NUV. FUV exposure times varied from 732 to 2518 seconds, while NUV exposure times ranged from 70 to 2788 seconds. Each orbit’s visit structure included an acquisition (ACQ) image, with an acquisition time between 4 seconds and 40 seconds. One orbit was used for the brightest WD, while 2 to 3 orbits were required for fainter, higher-extinction sources.
After the variability selection described in Section 2.1, 19 out of our 32 faint standards have STIS UV spectra to complement their optical ground-based spectra. The specific WDs that had this additional coverage are indicated in Table 1 and Table 3. Further information on the STIS UV spectra can be found in the preliminary analysis by Bohlin et al. (2024).
@@ 3 The Statistical Model
In this section we define the DAmodel statistical model, building on techniques previously developed by Narayan et al. (2019) and Axelrod et al. (2023). In Section 3.1 we describe the forward model to construct our dusty SEDs. In Section 3.2 we describe how we use the dusty SEDs in our likelihoods to model the HST/WFC photometry and instrumental systematics. In Section 3.3 we describe how we adapt the dusty SEDs to model observed optical and STIS UV spectra. In Section 3.4 we describe how we combine the different elements of the statistical model to make a fully hierarchical Bayesian model, which is also illustrated in Figure 3. Finally, in Section 3.5 we define the choice of non-informative priors on each inferred parameter.
@@ 3.1 Synthetic Spectral Energy Distribution
[…]
The Hubeny grid is defined from 900 Å to 320,000 Å with resolution of $R=$100,000 for DA WDs in the range of 15,000 K < $T_{\text{eff}}$ < 70,000 K and 7 < $\log\text{g}$ < 9.5. Our model’s unreddened SED $\mathcal{F}(\lambda;\log g,T_{\text{eff}})$ is determined by the cubic interpolation of the theoretical grid. The SED is then reddened according to the Gordon et al. (2023) dust relation where the extinction $A(\lambda;A_{V},R_{V})$ is applied such that
@@ 3.2 Photometric Modelling
$P\big{(}\hat{m}^{X}_{i,c,s}|\ \bm{\Theta^{s}_{\text{WD}}},\mu_{s},\bm{\theta_{\text{phot}}^{X}},\Delta Z_{X}^{c}\big{)}=\\ T\big{(}\hat{m}^{X}_{i,c,s}|\ m^{X}_{s}-Z_{X}-\Delta Z_{X}^{c}-\alpha_{X}(m^{X}_{s}-15),\sqrt{\sigma^{X\hskip 8.5359pt2}_{s,i,c}+\sigma_{\text{int}}^{X\hskip 2.84544pt2}},\nu^{X}\big{)}$ (4)
where $T(x|\ \bar{\mu},\sigma,\nu)$ is a student-T distribution with mean $\bar{\mu}$, scale $\sigma$ and $\nu$ degrees of freedom. We choose a student-T distribution to model the photometry due to its broader tails, allowing flexibility and robustness in modelling outliers. The measurement error for each individual photometric observation $i$ is represented by $\sigma^{X}_{s,i,c}$. Each HST band $X$ has its own set of parameters $\bm{\theta_{\text{phot}}^{X}}=(Z_{X},\sigma_{\text{int}}^{X},\nu^{X},\alpha_{X})^{T}$ containing a zeropoint $Z_{X}$, excess dispersion $\sigma_{\text{int}}^{X}$, degree of freedom $\nu^{X}$ and CRNL coefficient $\alpha_{X}$ (Riess et al., 2019; Bohlin & Deustua, 2019; Axelrod et al., 2023). The excess dispersion parameter $\sigma_{\text{int}}^{X}$, first introduced in Narayan et al. (2019), accounts for any additional band-dependent systematic scatter that is not accounted for in the measurement error of each observed object. If there are no additional systematics and the measurement errors describe the data perfectly, the model will infer $\sigma_{\text{int}}^{X}$ to be zero. We only consider a CRNL coefficient in the F160W band and set the coefficients for the remaining bands to zero. There is also an additional zeropoint offset $\Delta Z_{X}^{c}$ accounting for changes in zeropoints between Hubble cycles (Narayan et al., 2019; Calamida et al., 2022a; Marinelli et al., 2024) and differences in photometric settings. For at least one Hubble Cycle we must set $\Delta Z_{X}^{c}=0$ to avoid degeneracy with $Z_{X}$. For our analysis we set $\Delta Z_{X}^{25}=0$, then infer $\Delta Z_{X}^{20}$, and $\Delta Z_{X}^{22}$.
@@
$P\big{(}\hat{m}^{X}_{i,c,p}|\ \bm{\theta_{\text{phot}}^{X}},\Delta Z_{X}^{c}\big{)}=\\ T\big{(}\hat{m}^{X}_{i,c,p}|\ \bar{m}^{X}_{p}-Z_{X}-\Delta Z_{X}^{c}-\alpha_{X}(\bar{m}^{X}_{p}-15),\sqrt{\sigma^{X\hskip 8.5359pt2}_{s,i,c}+\sigma_{\text{int}}^{X\hskip 2.84544pt2}},\nu^{X}\big{)}$ (5)
@@ 3.3 Spectroscopic Modelling
$f\big{(}\lambda;\bm{\Theta^{s}_{\text{WD}}},\gamma^{j}_{s},\text{FWHM}_{s}^{j}\big{)}=\\ \frac{\tilde{F}(\lambda;\bm{\Theta^{s}_{\text{WD}}})}{4\pi\gamma_{j,s}^{2}}\hskip 2.84544pt\ast\hskip 2.84544ptN\big{(}\lambda|\ 0,\sigma_{R}(\text{FWHM}_{s}^{j})^{2}\big{)}$ (6)
The normalisation factor $\gamma^{j}_{s}$ acts as a nuisance factor to scale the synthetic spectrum to match observed spectrum $j$. The Gaussian kernel was introduced in Narayan et al. (2019) to smooth the synthetic spectrum to be consistent with the observed spectrum’s resolution, where absorption features are less sharp. The standard deviation of the kernel is defined as
@@
where $R$ is the resolution of the spectrum in Å per spectral unit and FWHM determines the extent of the smoothing.
@@
We choose to place the $M=10$ spline knots equally between the minimum and maximum wavelength coverage of each observed spectrum $j$. The spline knot locations avoid absorption features and are far enough apart to no add to them. We add the spline to the smoothed synthetic SED to make the modelled synthetic spectrum $S(\lambda;\bm{\Theta^{s}_{\text{WD}}},\bm{\theta}^{j,s}_{\text{spec}})$ using
@@
$P(\bm{\hat{S}}^{j}_{s}|\ \bm{\Theta^{s}_{\text{WD}}},\bm{\theta}^{j,s}_{\text{spec}})=\prod_{k=1}^{K_{s}^{j}}N\bigg{(}\hat{S}^{k,j}_{s}\bigg{|}\ S\big{(}\lambda^{s,j}_{k},\bm{\Theta^{s}_{\text{WD}}},\bm{\theta}^{j,s}_{\text{spec}}\big{)},\sigma^{j,s\hskip 1.42271pt2}_{k}\bigg{)}$ (10)
@@ 3.4 Hierarchical Model
For sampling the hierarchical model we use Hamiltonian Monte Carlo (Duane et al., 1987), a technique that uses gradient information (something that neural network architectures also use) for faster convergence on the posterior. We use the No-U-Turn Sampler (NUTS; Hoffman & Gelman 2011) in numpyro (Bingham et al., 2019; Phan et al., 2019), a probabilistic programming package for automatic differentiation and just-in-time (JIT) compilation built with JAX (Bradbury et al., 2018). Using this software combination on a GPU node, the implementation takes approximately 30 minutes in total to produce four chains of 500 burn-in samples and 500 posterior samples, amounting to an average effective sample size of 1277. This is a significant speed-up compared to the previous implementation of the non-hierarchical model from Narayan et al. (2019) that would take roughly one CPU day per object to converge. When assessing the quality of our chains we diagnose convergence by confirming the Gelman-Rubin $\hat{R}$ statistic is less than 1.02 for each of the 3653 inferred parameters (Gelman & Rubin, 1992; Vehtari et al., 2021). We also find that none of the chains result in divergent transitions (Betancourt et al., 2014; Betancourt, 2017). The implementation of DAmodel can be found at: https://github.com/benboyd97/DAmodel.
@@ 3.5 Priors
In our modelling we use non-informative priors $P(\bm{\Theta})$ on our parameters. The priors on intrinsic WD and dust parameters are defined as
$P(\log g^{s})$ $=U(\log g^{s}|\ 7\text{},9.5\text{ })$ (14)
$P(T_{\text{eff}}^{s})$ $=U(T_{\text{eff}}^{s}|\ 15000\text{ K },70000\text{ K })$ (15)
$P(R_{V}^{S}|\ \mu_{RV},\sigma_{RV})$ $=TN(R_{V}^{s}|\ \mu_{RV},\sigma_{RV}^{2},1.2,\infty)$ (16)
@@
where $HC(0,1)$ is a Half-Cauchy distribution with a scale of 1. The lower bound on $R_{V}$ of 1.2 is motivated by the Rayleigh scattering limit (Draine, 2003).
@@
$P(\sigma_{\text{int}}^{X})$ $=HC(\sigma_{\text{int}}^{X}|\ 0,1)$ (22)
$P(Z_{X})$ $=N(Z_{X}|\ \mu_{Z}^{X},1^{2})$ (23)
@@
$P(\text{FWHM}^{j}_{s})$ $=TN(\text{FWHM}^{j}_{s}|\ 0,8^{2},0,50)$ (28)

### v2

@@ 2.3 Ground-Based Optical Spectra
Ground-based optical spectra were obtained with the Gemini North and South telescopes, the Multiple Mirror Telescope (MMT) and Southern Astrophysical Research (SOAR) Telescope, as described in Calamida et al. (2019) and Calamida et al. (2022b). The observations of the standards with Gemini spectrographs ($1.0^{\prime\prime}$ to $1.5^{\prime\prime}$ slit) cover a continuous range from 3500 Å to 6360 Å. The MMT observations of the standards ($1.0^{\prime\prime}$ to $1.25^{\prime\prime}$ slit) cover a wavelength range from 3400 Å to 8400 Å. The DA WDs were observed with SOAR ($1.07^{\prime\prime}$ slit) in a wavelength range between 3850 Å and 7100 Å. The spectrographs did not operate on a linear pixel-to-angstrom scale, so the data was resampled onto a uniform linear grid after extraction, with a sampling of 1 Å/pixel for Gemini, 2 Å/pixel for MMT and 2 Å/pixel for SOAR. For objects with multiple spectroscopic observations in the optical, we default to using the same spectrum used in the Narayan et al. (2019) and Axelrod et al. (2023) analyses.
@@ 2.4 Space-Based STIS UV Spectra
To better characterise and ultimately correct for dust effects, the spectroscopic observations were extended into the UV regime. The HST Space Telescope Imaging Spectrograph (STIS) was used to observe 19 of the faint DA WDs in program 16764 (PI Gautham Narayan). These stars had apparent magnitudes in the F275W filter between 15 mag and 18 mag. Spectra were composed of observations using both FUV and NUV detectors with the G140L and G230L gratings, covering a combined wavelength range of 1150 Å to 3180 Å. The resolution was resampled to 1 Å/pixel for our analysis. The specific WDs that had this additional coverage are indicated in Table 1 and Table 3. Further information on the STIS UV spectra can be found in the preliminary analysis by Bohlin et al. (2025).
@@ 3 The Statistical Model
In this section we define the DAmodel statistical model, building on techniques previously developed by Narayan et al. (2019) and Axelrod et al. (2023). In Section 3.1 we describe the forward model to construct our dusty SEDs. In Section 3.2 we describe how we use the dusty SEDs in our likelihoods to model the HST/WFC photometry and instrumental systematics. In Section 3.3 we describe how we adapt the dusty SEDs to model observed optical and STIS UV spectra. In Section 3.4 we describe how we combine the different elements of the statistical model to make a fully hierarchical Bayesian model, which is also illustrated in Figure 3. Finally, in Section 3.5 we define the choice of priors on each inferred parameter.
@@ 3.1 Synthetic Spectral Energy Distribution
Narayan et al. (2019) and Axelrod et al. (2023) previously used a non-local thermodynamic equilibrium (NLTE) model grid computed by Tlusty v202 code (Hubeny & Lanz, 2017). In this work we use the recently updated NLTE grid [^1] The updated theoretical model spectra for hot DA white dwarfs used in this work can be found at: www.as.arizona.edu/~hubeny/pub/DAgrid24.tar.gz computed with Tlusty v208 (Hubeny et al., 2021) that provides us with improved modelling of Lyman-$\alpha$ at UV wavelengths, necessary for analysis of the STIS spectra. There are still future updates to come with the theoretical DA WD grids, particularly with development of hydrogen Lyman lines Stark broadening, as well as improvements in the UV for cooler DA WDs at $T_{\text{eff}}$ < 20 000 K. Our method is general and can seamlessly integrate future grids as they become available.
The new Hubeny grid is defined from 900 Å to 32 $\mu$m with resolution of $R=$100 000 for DA WDs in the range of 15 000 K < $T_{\text{eff}}$ < 70 000 K and 7 < $\log\text{g}$ < 9.5. The resolution of the grid is 0.25 for $\log\text{g}$ and 250 K, 500 K and 1000 K for effective temperatures up to 20 000 K, 40 000 K and 70 000 K respectively. Our model’s unreddened SED $\mathcal{F}(\lambda;\log g,T_{\text{eff}})$ is determined by the cubic interpolation of the theoretical grid. The SED is then reddened according to the Gordon et al. (2023) dust relation (Gordon et al., 2009, 2021; Fitzpatrick et al., 2019; Decleir et al., 2022; Gordon, 2024a, b) where the extinction $A(\lambda;A_{V},R_{V})$ is applied such that
@@ 3.2 Photometric Modelling
$P\big{(}\hat{m}^{X}_{i,c,s}|\ \bm{\Theta^{s}_{\text{WD}}},\mu_{s},\bm{\theta_{\text{phot}}^{X}},\Delta Z_{X}^{c}\big{)}=\\ \mathcal{T}\big{(}\hat{m}^{X}_{i,c,s}|\ m^{X}_{s}-Z_{X}-\Delta Z_{X}^{c}-\alpha_{X}(m^{X}_{s}-15),\sqrt{\sigma^{X\hskip 8.5359pt2}_{i,c,s}+\sigma_{\text{int}}^{X\hskip 2.84544pt2}},\nu^{X}\big{)}$ (4)
where $\mathcal{T}(x|\ \bar{\mu},\sigma,\nu)$ is a student-T distribution with mean $\bar{\mu}$, scale $\sigma$ and $\nu$ degrees of freedom. We choose a student-T distribution to model the photometry due to its broader tails, allowing flexibility and robustness in modelling outliers. The measurement error for each individual photometric observation $i$ is represented by $\sigma^{X}_{i,c,s}$. Each HST band $X$ has its own set of parameters $\bm{\theta_{\text{phot}}^{X}}=(Z_{X},\sigma_{\text{int}}^{X},\nu^{X},\alpha_{X})^{T}$ containing a zeropoint $Z_{X}$, excess dispersion $\sigma_{\text{int}}^{X}$, degree of freedom $\nu^{X}$ and count-rate nonlinearity (CRNL) coefficient $\alpha_{X}$ (Riess et al., 2019; Bohlin & Deustua, 2019; Axelrod et al., 2023). The excess dispersion parameter $\sigma_{\text{int}}^{X}$ accounts for any additional band-dependent systematic scatter that is not accounted for in the measurement error of each observed object. If there are no additional systematics and the measurement errors describe the data perfectly, the model will infer $\sigma_{\text{int}}^{X}$ to be zero. We only consider a CRNL coefficient in the F160W band and set the coefficients for the remaining bands to zero. There is also an additional zeropoint offset $\Delta Z_{X}^{c}$ accounting for changes in zeropoints between Hubble cycles (Narayan et al., 2019; Calamida et al., 2022a; Marinelli et al., 2024) and differences in photometric settings. For at least one Hubble Cycle we must set $\Delta Z_{X}^{c}=0$ to avoid degeneracy with $Z_{X}$. For our analysis we set $\Delta Z_{X}^{25}=0$, then infer $\Delta Z_{X}^{20}$, and $\Delta Z_{X}^{22}$.
@@
$P\big{(}\hat{m}^{X}_{i,c,p}|\ \bm{\theta_{\text{phot}}^{X}},\Delta Z_{X}^{c}\big{)}=\\ \mathcal{T}\big{(}\hat{m}^{X}_{i,c,p}|\ \bar{m}^{X}_{p}-Z_{X}-\Delta Z_{X}^{c}-\alpha_{X}(\bar{m}^{X}_{p}-15),\sqrt{\sigma^{X\hskip 8.5359pt2}_{i,c,p}+\sigma_{\text{int}}^{X\hskip 2.84544pt2}},\nu^{X}\big{)}$ (5)
@@ 3.3 Spectroscopic Modelling
$f\big{(}\lambda;\bm{\Theta^{s}_{\text{WD}}},\gamma^{j}_{s},\text{FWHM}_{s}^{j}\big{)}=\\ \frac{\tilde{F}(\lambda;\bm{\Theta^{s}_{\text{WD}}})}{4\pi\gamma_{j,s}^{2}}\hskip 2.84544pt\ast\hskip 2.84544pt\mathcal{N}\big{(}\lambda|\ 0,\sigma_{R}(\text{FWHM}_{s}^{j})^{2}\big{)}$ (6)
[…]
@@
where $R$ is the resolution of the spectrum in Å per spectral unit and FWHM determines the extent of the smoothing. We adopt this simple and consistent smoothing approach because we are combining spectra from different ground and space-based instruments, each operating at different wavelengths with varying line-spread functions.
@@
We choose to place the $M=10$ spline knots equally between the minimum and maximum wavelength coverage of each observed spectrum $j$. The spline knot locations avoid absorption features and are far enough apart to not artificially add to their depths. We add the spline to the smoothed synthetic SED to make the modelled synthetic spectrum $S(\lambda;\bm{\Theta^{s}_{\text{WD}}},\bm{\theta}^{j,s}_{\text{spec}})$ using
@@
$P(\bm{\hat{S}}^{j}_{s}|\ \bm{\Theta^{s}_{\text{WD}}},\bm{\theta}^{j,s}_{\text{spec}})=\prod_{k=1}^{K_{s}^{j}}\mathcal{N}\bigg{(}\hat{S}^{k,j}_{s}\bigg{|}\ S\big{(}\lambda^{s,j}_{k};\bm{\Theta^{s}_{\text{WD}}},\bm{\theta}^{j,s}_{\text{spec}}\big{)},\sigma^{j,s\hskip 1.42271pt2}_{k}\bigg{)}$ (10)
@@ 3.4 Hierarchical Model
For sampling the hierarchical model we use Hamiltonian Monte Carlo (Duane et al., 1987), a technique that uses gradient information (something that neural network architectures also use) for faster convergence on the posterior. We use the No-U-Turn Sampler (NUTS; Hoffman & Gelman 2014) in numpyro (Bingham et al., 2019; Phan et al., 2019), a probabilistic programming package for automatic differentiation and just-in-time (JIT) compilation built with JAX (Bradbury et al., 2018). Using this software combination on a GPU node, the implementation takes approximately 30 minutes in total to produce four chains of 500 burn-in samples and 500 posterior samples, amounting to an average effective sample size of 1277. This is a significant speed-up compared to the previous implementation of the non-hierarchical model from Narayan et al. (2019) that would take roughly one CPU day per object to converge. When assessing the quality of our chains we diagnose convergence by confirming the Gelman-Rubin $\hat{R}$ statistic is less than 1.02 for each of the 3653 inferred parameters (Gelman & Rubin, 1992; Vehtari et al., 2021). We also find that none of the chains result in divergent transitions (Betancourt et al., 2014; Betancourt, 2017). The implementation of DAmodel can be found at: https://github.com/benboyd97/DAmodel.
@@ 3.5 Priors
In our modelling we use priors $P(\bm{\Theta})$ on our parameters. The priors on intrinsic WD and dust parameters are defined as
$P(\log g^{s})$ $=\mathcal{U}(\log g^{s}|\ 7\text{},9.5\text{ })$ (14)
$P(T_{\text{eff}}^{s})$ $=\mathcal{U}(T_{\text{eff}}^{s}|\ 15000\text{ K },70000\text{ K })$ (15)
$P(R_{V}^{S}|\ \mu_{RV},\sigma_{RV})$ $=\mathcal{TN}(R_{V}^{s}|\ \mu_{RV},\sigma_{RV}^{2},1.2,\infty)$ (16)
@@
where $\mathcal{HC}(0,1)$ is a Half-Cauchy distribution with a scale of 1. The lower bound on $R_{V}$ of 1.2 is motivated by the Rayleigh scattering limit (Draine, 2003).
@@
$P(\sigma_{\text{int}}^{X})$ $=\mathcal{HC}(\sigma_{\text{int}}^{X}|\ 0,1)$ (22)
$P(Z_{X})$ $=\mathcal{N}(Z_{X}|\ \mu_{Z}^{X},1^{2})$ (23)
@@
$P(\text{FWHM}^{j}_{s})$ $=\mathcal{TN}(\text{FWHM}^{j}_{s}|\ 0,8^{2},0,50)$ (28)